# Preliminary working paper

### v1

@@ 1 Introduction
In this paper I explore a new mechanism that contributes to the explanation on why human-AI collaboration, particularly in AI recommendation systems, often falls short of expectations. This paper focuses on studying the impact of non-instrumental image concerns on AI utilization rates. Are workers worried about how others perceive them when they use AI? If so, image concerns may discourage workers from making decisions based solely on accuracy. For instance, a worker who consistently follows AI recommendations may appear to exert low effort or contribute little value (Weitzner 2024), while one who rarely follows them could seem resistant to technological adaptation. Although image concerns may affect payoffs directly or indirectly, I focus here on showing that they can also reflect psychological forces that operate independently of monetary incentives.
@@ 1.1 Related Literature
High-stakes professional environments have faced the perils and challenges of implementing AI recommendations. Medical clinics (Agarwal et al. 2023), judicial courts (Angelova et al. 2024), and HR offices (Hoffman et al. 2018) are just a few examples of settings where the benefits of accessible AI recommendations have yet to materialize. A common denominator across these studies is the systematic under-utilization of AI recommendations, a phenomenon previously coined as algorithmic aversion (Dietvorst et al. 2015). The most common explanation for these underwhelming outcomes centers on overconfidence and other deviations from Bayesian updating. Using controlled experiment environments, Caplin et al. (2024) provide evidence that calibration and belief formation play a key role in determining the gains from working with AI recommendations, while Serra-Garcia and Gneezy (2023) show that the timing of algorithmic feedback matters: receiving feedback before forming one’s own beliefs significantly increases adoption. If these explanations alone were sufficient to account for the issue, then assuming that professionals only care about maximizing accuracy would be reasonable. However, there is already evidence suggesting that working with AI may influence the preferences of the parties involved. [^2] Specifically in the context of recommendation systems, McLaughlin and Spiess (2024) propose a model of reference-dependent utility in which the reference point is influenced by the AI recommendation. Albright (2024) provides empirical evidence suggesting that algorithmic recommendations may offer judges reputational cover when making visible mistakes. In an AI oversight setting, Almog et al. (2025a) show that being overruled by Hawk-Eye in public asymmetrically altered the costs of different error types for professional tennis umpires. Grimon and Mills (2025) provide one of the few bright spots in the economics literature, documenting collaboration gains when humans operate under AI recommendations. Their study examines a setting in which Child Protective Services workers had to decide whether to send an investigator to screen a reported family. A key implementation feature was that workers were encouraged by their managers to consult the AI tool during team meetings and received training to understand their comparative advantage relative to the algorithm. [^3] Workers were aware that the algorithm was trained only on family background information, while details from the call report were treated as “private information” available only to the worker. These distinctions in the implementation may have helped reduce the stigma associated with using the tool, an important factor that may not have been addressed in other settings where human-AI collaboration failed to generate complementarities. I present evidence that non-instrumental image concerns may help reconcile previously mixed empirical findings and offer guidance on the environments under which human–AI collaboration yields better outcomes.
@@
Image concerns can influence consequential personal decisions, including voting (Dellavigna et al. 2017), credit consumption (Bursztyn et al. 2018), educational investment (Bursztyn et al. 2019), political contributions (Perez-Truglia and Cruces 2017), and childhood immunization (Karing 2024). This study adds to the growing body of research on image concerns (for a review, see Bursztyn and Jensen (2017)) by identifying a new domain in which these forces shape behavior: human–AI collaboration. Image concerns are also deeply relevant in the workplace, where non-monetary incentives play a central role. Workers often derive utility not just from wages, but also from meaning (Cassar and Meier 2018), pride, respect (Ellingsen and Johannesson 2007, 2008), and how they are compared to fellow workers (Mas and Moretti 2009). This paper’s findings that non-instrumental image concerns can deter AI utilization support the view that non-monetary aspects of the workplace can meaningfully affect workers’ decisions.
@@ 2.5 Design Choices
Video call treatment. A video call provides the most credible intervention to genuinely convey that another person is observing participants’ AI usage; other alternatives, such as a personalized message, are less persuasive given how easily they can be automated today. A face-to-face interaction provides the best chance to experimentally induce image concerns from a one-time interaction with a stranger. Another option was to host the experiment in person at the university lab. However, it is difficult to ignore the student–experimenter connection in that setting, as students who participate in lab studies often do so repeatedly and may develop an ongoing relationship with the lab.
@@ 3.1 Treatment Effects
Since the option to revise an answer was only offered when the AI recommendation differed from the participant’s initial choice, a valid concern is that the treatment may have influenced behavior at the initial choice stage. Specifically, the treatment could have prompted participants to exert more effort initially, leading them to encounter a systematically different set of images for which the AI recommendation would later be shown. Having this in mind, I deliberately selected a task with minimal learning potential and limited returns to additional effort, aiming to ensure that the revision stage of the decision process remained as comparable as possible across treatment conditions. Regressing initial correct answers on Response time (controlling for image-level noise and participant fixed effects), the estimated effect is a 2.2 p.p. decrease in accuracy per additional 10 seconds spent. The effect is not only small considering the average response time was 14 seconds, but also negative. This indicates that higher performance in this task tends to come from faster decisions, reinforcing the idea that effort offers limited returns. Finally, Columns 3 and 5 of Table 2 provide evidence against a distortion in initial choice behavior, showing no significant treatment effects on either accuracy or response time at the initial choice stage.
@@ 4 Conclusion
Importantly, the effects of image concerns are not the same for everyone. The treatment had a stronger impact on minority participants, suggesting that even seemingly neutral AI recommendation systems can unintentionally reinforce inequality when social dynamics are overlooked. Much of the algorithmic fairness literature has focused on adjusting algorithm outputs (Liang et al. 2025). However, these results highlight the importance of understanding how people interact with algorithms, and how social background can shape those interactions. This points to the need for a broader approach to algorithmic fairness, one that considers not just the algorithm’s output, but the final outcomes that emerge from human–AI collaboration.

### v2

@@ 1 Introduction
In this paper I explore a new mechanism that contributes to the explanation on why human-AI collaboration, particularly in AI recommendation systems, often falls short of expectations. This paper focuses on studying the impact of non-instrumental image concerns (also referred to as hedonic motives by Bursztyn and Jensen (2017)) on AI utilization rates. Are workers worried about how others perceive them when they use AI? If so, image concerns may discourage workers from making decisions based solely on accuracy. For instance, a worker who consistently follows AI recommendations may be perceived as exerting low effort or contributing little value (Weitzner 2024), while one who rarely follows them may appear resistant to technological adaptation. These examples illustrate that image concerns could plausibly push AI utilization in either direction. However, recent evidence suggests that social norms are beginning to emerge against the use of AI. For example, Ling and Imas (2025) find that social desirability bias leads people to under-report their AI usage. Similarly, Awuah et al. (2025) show that human evaluators tend to penalize LLM-generated job applications, even though this is not formally specified as an evaluation criterion. While image concerns may affect payoffs either directly or indirectly, my focus here is to show that they can also reflect psychological forces that operate independently of monetary incentives.
@@ 1.1 Related Literature
High-stakes professional environments have faced the perils and challenges of implementing AI recommendations. Medical clinics (Agarwal et al. 2023), judicial courts (Angelova et al. 2024), and HR offices (Hoffman et al. 2018) are just a few examples of settings where the benefits of accessible AI recommendations have yet to materialize. A common denominator across these studies is the systematic under-utilization of AI recommendations, a phenomenon previously coined as algorithmic aversion (Dietvorst et al. 2015). The most common explanation for these underwhelming outcomes centers on overconfidence and other deviations from Bayesian updating. Using controlled experiment environments, Caplin et al. (2024) provide evidence that calibration and belief formation play a key role in determining the gains from working with AI recommendations, while Serra-Garcia and Gneezy (2023) show that the timing of algorithmic feedback matters: receiving feedback before forming one’s own beliefs significantly increases adoption. If belief formation explanations alone were sufficient to account for the issue, it would be reasonable to assume that professionals only care about maximizing accuracy. However, there is already evidence suggesting that working with AI may influence the preferences of the parties involved. [^2] Specifically in the context of recommendation systems, McLaughlin and Spiess (2024) propose a model of reference-dependent utility in which the reference point is influenced by the AI recommendation. Albright (2024) provides empirical evidence suggesting that algorithmic recommendations may offer judges reputational cover when making visible mistakes. In an AI oversight setting, Almog et al. (2025a) show that being overruled by Hawk-Eye in public asymmetrically altered the costs of different error types for professional tennis umpires. Grimon and Mills (2025) provide one of the few bright spots in the economics literature, documenting collaboration gains when humans operate under AI recommendations. Their study examines a setting in which Child Protective Services workers had to decide whether to send an investigator to screen a reported family. A key implementation feature was that workers were encouraged by their managers to consult the AI tool during team meetings and received training to understand their comparative advantage relative to the algorithm. [^3] Workers were aware that the algorithm was trained only on family background information, while details from the call report were treated as “private information” available only to the worker. These distinctions in the implementation may have helped reduce the stigma associated with using the tool, an important factor that may not have been addressed in other settings where human-AI collaboration failed to generate complementarities. I present evidence that non-instrumental image concerns may help reconcile previously mixed empirical findings and offer guidance on the environments under which human–AI collaboration yields better outcomes.
@@
Image concerns can influence consequential personal decisions, including voting (Dellavigna et al. 2017), credit consumption (Bursztyn et al. 2018), educational investment (Bursztyn et al. 2019), political contributions (Perez-Truglia and Cruces 2017), and childhood immunization (Karing 2024). This study adds to the growing body of research on image concerns (for a review, see Bursztyn and Jensen (2017)) by identifying a new domain in which these forces shape behavior: human–AI collaboration. Image concerns are also deeply relevant in the workplace, where non-monetary incentives play a central role. Workers often derive utility not just from wages, but also from meaning (Cassar and Meier 2018; Ashraf et al. 2025), pride, respect (Ellingsen and Johannesson 2007, 2008), and how they are compared to fellow workers (Mas and Moretti 2009). This paper’s findings that non-instrumental image concerns can deter AI utilization support the view that non-monetary aspects of the workplace can meaningfully affect workers’ decisions.
@@ 2.5 Design Choices
Video call treatment. A video call provides the most credible intervention to genuinely convey that another person is observing participants’ AI usage; other alternatives, such as a personalized message, are less persuasive given how easily they can be automated today. Face-to-face interactions provide the strongest opportunity to experimentally induce image concerns in a one-time encounter with a stranger. This method has been widely used in experimental research to isolate the non-instrumental dimension of image concerns, demonstrating that interacting with a stranger can influence behaviors such as voting (Dellavigna et al. 2017) and charitable giving (Dellavigna et al. 2012). Another option was to host the experiment in person at the university lab. However, it is difficult to ignore the student–experimenter connection in that setting, as students who participate in lab studies often do so repeatedly and may develop an ongoing relationship with the lab.
@@ 3.1 Treatment Effects
Since the option to revise an answer was only offered when the AI recommendation differed from the participant’s initial choice, a valid concern is that the treatment may have influenced behavior at the initial choice stage. Specifically, the treatment could have prompted participants to exert more effort initially, leading them to encounter a systematically different set of images for which the AI recommendation would later be shown. [^10] Agarwal et al. (2025) Illustrate how effort responses should be considered when designing human-AI collaboration systems. Having this in mind, I deliberately selected a task with minimal learning potential and limited returns to additional effort, aiming to ensure that the revision stage of the decision process remained as comparable as possible across treatment conditions. Regressing initial correct answers on Response time (controlling for image-level noise and participant fixed effects), the estimated effect is a 2.2 p.p. decrease in accuracy per additional 10 seconds spent. The effect is not only small considering the average response time was 14 seconds, but also negative. This indicates that higher performance in this task tends to come from faster decisions, reinforcing the idea that effort offers limited returns. Finally, Columns 3 and 5 of Table 2 provide evidence against a distortion in initial choice behavior, showing no significant treatment effects on either accuracy or response time at the initial choice stage.
@@ 4 Conclusion
Importantly, the effects of image concerns are not the same for everyone. The treatment had a stronger impact on minority participants, suggesting that even seemingly neutral AI recommendation systems can unintentionally reinforce inequality when social dynamics are overlooked. Much of the algorithmic fairness literature has focused on adjusting algorithm outputs (Liang et al. 2025). However, these results highlight the importance of understanding how people interact with algorithms, and how social background can shape those interactions. This extends the call for a broader approach to algorithmic fairness (Ge et al. 2025), one that considers not just the algorithm’s output, but the final outcomes that emerge from human–AI collaboration.